Fill in the Action and Goto tables for an LALR parser for the grammar given below, where the start symbol is S. Although the grammar is ambiguous, your parser must not be ambiguous: an ambiguous string, such as iiaea, must be parsed in the “correct” manner, using the usual disambiguating rule for the “dangling else” problem.

1. S ⟶ a_{2}

2. S ⟶ w_{3}S_{4}

3. S ⟶ i_{5}S_{6}

4. S ⟶ i_{5}S_{6}e_{7}S_{8}

5. S ⟶ b_{9}L_{10}n_{11}

6. L ⟶ L_{10}S_{12}

7. L ⟶ ϵ